# The Learning Dynamics of Subword Segmentation for Morphologically Diverse Languages

### v1

@@ 8.1 Computational complexity
Like other models that marginalise over multiple tokenisations (kreutzer-sokolov-2018-learning; he-etal-2020-dynamic; meyer-buys-2022-subword), T-SSLM is computationally more expensive than standard LM training. Despite the dynamic program of Eq. 3, marginalisation adds unavoidable cost by computing probabilities for multiple candidate subword sequences (standard LMs compute the likelihood for a single subword sequence). In practice, T-SSLM training times increase by an order of magnitude ($\times 10$) over tokenisation-based LMs.
The added computational complexity of learning subwords prevents us from pretraining on larger datasets. It is an open question whether our findings would scale – whether similar subword learning trajectories would be observed for larger-scale pretraining and higher-resourced languages. This would require pretraining a large-scale SSLM from scratch on a massive corpus, which is not possible given the compute resources at our disposal. The claims of this paper are limited to smaller-scale pretraining of low-resource languages. This is a relevant and important use case, since data scarcity and computational constraints often co-occur in practice (ahia-etal-2021-low-resource).
@@ 8.2 Linguistic scope
(a) Setswana
(b) English
(c) isiXhosa
Figure 4: Fertility distributions (subwords per segmented word) across the four training stages identified in this study. All languages exhibit increasing fertility as pretraining progresses, with a more pronounced shift during finetuning. Changes are especially dramatic for isiXhosa, whose complex morphology leads to larger distributional shifts than Setswana or English.
(a) Setswana-pretrained
(b) English-pretrained
Figure 5: Boundary overlap between isiXhosa morphemes and subwords learned by our T-SSLMs pretrained on Setswana (left) and English (right). During isiXhosa finetuning, the models adjusts their subword segmentation to align with the morphological boundaries of the new language.
(a) Full prompt template examples for finetuning
South Africa + leaderName + Cyril Ramaphosa # # =uCyril Ramaphosa yinkokheli yomzantsi Afrika.
Ethiopia + currency + Ethiopian birr # # =Imali yase-Ethiopia yi-Ethiopian Birr.
Denmark + capital + Copenhagen # # =ICopenhagen likomkhulu laseDenmark.
(b) Incomplete prompt template examples for testing
Romania + capital + Bucharest # # =
India + currency + Indian rupee # # =
Netherlands + leaderName + Mark Rutte # # =
Table 6: To finetune decoder-only PLMS for T2X isiXhosa data-to-text generation, we cast the task as prompt completion. Here we demonstrate the prompt template, “{input text} # # ={output text}”, with examples from the T2X dataset. During finetuning (a), we only maximise the generation of text after the equals sign “=”. During testing (b), the special tokens “# # =” prompt output generation.
@@ A.1 Mixture model
T-SSLM encodes the preceding text as an untokenised character sequence, using a character-level Transformer. This discards information about segmentation history, but achieves computationally feasible conditioning. For next-subword prediction, we adapt the approach of meyer-buys-2022-subword to use a Transformer encoding. To condition next-subword prediction on preceding text, we pass the final output embedding to two decoding subnetworks, a character-level model and a softmax layer over a subword lexicon.
@@ A.2 Finetuning algorithm
While we can use the existing dynamic programming algorithm of meyer-buys-2022-subword for pretraining, we have to adapt it to finetune our T-SSLM for prompt-based text generation. Their dynamic program is represented by Eq. 3, which computes (and thereby maximises) the likelihood for a full text sequence. This is not compatible with open-ended text generation, which are cast as text completion tasks: given an input prompt $C$ as context, generate the expected output $O$. The model is finetuned to maximise $p(O|C)$ directly, instead of finetuning on generating both $C$ and $O$. Such completion-only finetuning is standard practice in prompt-based generation.
@@ A.3 Decoding
Standard beam search operates on one vocabulary, but the mixture model of T-SSLM combines two vocabularies, subwords and characters, complicating text generation. We adapt dynamic decoding, previously proposed for encoder-decoder subword segmental models, for decoder-only generation. Dynamic decoding (meyer-buys-2023-subword) generates the target translation one character at a time and predicts subword boundaries based on mixture model probabilities. We adapt this algorithm for our model by conditioning next-character probabilities on preceding prompt context, as opposed to a source-language sentence. During generation, we use a beam size of 5 in our dynamic decoding algorithm.
@@ Appendix C Baseline Hyperparameters
For BPE, ULM, and BPE-dropout, we use a subword vocabulary size match the subword lexicon of our SSLM (5k for Setswana and English, 10k for isiXhosa). We implement our baselines with the Huggingface Transformers library (wolf-etal-2020-transformers), training LMs of the same size as our isiXhosa SSLMs (6 layers, 8 attention heads, 512-dimensional embeddings, maximum sequence length of 512). We pretrain the baseline LMs using a batch size of 64 and the default Huggingface pretraining hyperparameters for causal language modelling: a learning rate of 5e-5, no warmup, a linear scheduler, and a dropout rate of 0.1.

### v2

@@ 8.1 Computational complexity
Like other models that marginalise over multiple tokenisations (Kreutzer and Sokolov, 2018; He et al., 2020; Meyer and Buys, 2022), T-SSLM is computationally more expensive than standard LM training. Despite the dynamic program of Eq. 3, marginalisation adds unavoidable cost by computing probabilities for multiple candidate subword sequences (standard LMs compute the likelihood for a single subword sequence). In practice, T-SSLM training times increase by an order of magnitude ($\times 10$) over tokenisation-based LMs.
The added computational complexity of learning subwords prevents us from pretraining on larger datasets. It is an open question whether our findings would scale – whether similar subword learning trajectories would be observed for larger-scale pretraining and higher-resourced languages. This would require pretraining a large-scale SSLM from scratch on a massive corpus, which is not possible given the compute resources at our disposal. The claims of this paper are limited to smaller-scale pretraining of low-resource languages. This is a relevant and important use case, since data scarcity and computational constraints often co-occur in practice (Ahia et al., 2021).
@@ A.1 Mixture model
T-SSLM encodes the preceding text as an untokenised character sequence, using a character-level Transformer. This discards information about segmentation history, but achieves computationally feasible conditioning. For next-subword prediction, we adapt the approach of Meyer and Buys (2022) to use a Transformer encoding. To condition next-subword prediction on preceding text, we pass the final output embedding to two decoding subnetworks, a character-level model and a softmax layer over a subword lexicon.
@@ A.2 Finetuning algorithm
While we can use the existing dynamic programming algorithm of Meyer and Buys (2022) for pretraining, we have to adapt it to finetune our T-SSLM for prompt-based text generation. Their dynamic program is represented by Eq. 3, which computes (and thereby maximises) the likelihood for a full text sequence. This is not compatible with open-ended text generation, which are cast as text completion tasks: given an input prompt $C$ as context, generate the expected output $O$. The model is finetuned to maximise $p(O|C)$ directly, instead of finetuning on generating both $C$ and $O$. Such completion-only finetuning is standard practice in prompt-based generation.
@@ A.3 Decoding
Standard beam search operates on one vocabulary, but the mixture model of T-SSLM combines two vocabularies, subwords and characters, complicating text generation. We adapt dynamic decoding, previously proposed for encoder-decoder subword segmental models, for decoder-only generation. Dynamic decoding (Meyer and Buys, 2023) generates the target translation one character at a time and predicts subword boundaries based on mixture model probabilities. We adapt this algorithm for our model by conditioning next-character probabilities on preceding prompt context, as opposed to a source-language sentence. During generation, we use a beam size of 5 in our dynamic decoding algorithm.
@@ Appendix C Baseline Hyperparameters
For BPE, ULM, and BPE-dropout, we use a subword vocabulary size match the subword lexicon of our SSLM (5k for Setswana and English, 10k for isiXhosa). We implement our baselines with the Huggingface Transformers library (Wolf et al., 2020), training LMs of the same size as our isiXhosa SSLMs (6 layers, 8 attention heads, 512-dimensional embeddings, maximum sequence length of 512). We pretrain the baseline LMs using a batch size of 64 and the default Huggingface pretraining hyperparameters for causal language modelling: a learning rate of 5e-5, no warmup, a linear scheduler, and a dropout rate of 0.1.